# DynoSAM: Open-Source Smoothing and Mapping Framework for Dynamic SLAM

### v2

@@ VII-A Metrics
$00$ $01$ $02$ $03$ $04$ $05$ $06$ $18$ $20$ L1 L2 S1 S2 avg. S4U
ATE($\mathrm{m}$) DynaSLAM II $1.29$ $2.31$ $0.91$ $\mathbf{0.69}$ $\mathbf{1.42}$ $\mathbf{1.34}$ $\mathbf{0.19}$ $\mathbf{1.09}$ $1.36$
VDO-SLAM $3.37$ $6.74$ $2.47$ $2.12$ $4.53$ $3.8$ $0.45$ $9.94$ $7.82$ $0.19$
ClusterSLAM - - - - - - - - - - - - - $0.53$
MVO $1.53$ $\mathbf{0.05}$
DynoSAM Pose $\mathbf{0.82}$ $\mathbf{2.00}$ $\mathbf{0.73}$ $0.82$ $2.01$ $1.58$ $0.31$ $1.84$ $\mathbf{1.26}$ $0.89$ $\mathbf{0.84}$ $\mathbf{0.25}$ $\mathbf{0.52}$ $0.63$ $0.11$
DynoSAM Motion $\mathbf{0.82}$ $\mathbf{2.00}$ $\mathbf{0.73}$ $0.82$ $2.01$ $1.58$ $0.31$ $1.84$ $\mathbf{1.26}$ $\mathbf{0.1}$ $\mathbf{0.84}$ $0.26$ $\mathbf{0.52}$ $\mathbf{0.51}$ $0.11$
[…]
VDO-SLAM $0.08$ $0.05$ $0.03$ $\mathbf{0.03}$ $0.06$ $0.03$ $0.1$ $0.03$ $0.04$ $0.77$
ClusterSLAM - - - - - - - - - - - - - $1.15$
MVO $0.19$ $0.76$
DynoSAM Pose $0.05$ $\mathbf{0.03}$ $\mathbf{0.02}$ $0.05$ $\mathbf{0.06}$ $0.06$ $0.05$ $0.04$ $\mathbf{0.04}$ $\mathbf{0.88}$ $\mathbf{0.69}$ $\mathbf{0.03}$ $\mathbf{0.30}$ $\mathbf{0.48}$ $0.69$
DynoSAM Motion $\mathbf{0.04}$ $\mathbf{0.03}$ $\mathbf{0.02}$ $0.05$ $\mathbf{0.06}$ $0.05$ $0.05$ $0.04$ $0.04$ $1.26$ $\mathbf{0.69}$ $\mathbf{0.03}$ $0.31$ $0.57$ $\mathbf{0.69}$
$\text{RPE}_{t}$($\mathrm{m}$) DynaSLAM II $\mathbf{0.04}$ $0.05$ $0.04$ $\mathbf{0.04}$ $0.07$ $0.06$ $0.02$ $0.05$ $0.07$
VDO-SLAM $0.09$ $0.15$ $0.05$ $0.09$ $0.14$ $0.11$ $0.04$ $0.09$ $0.30$ $0.12$
ClusterSLAM - - - - - - - - - - - - - $1.10$
MVO $0.07$ $\mathbf{0.004}$
DynoSAM Pose $\mathbf{0.04}$ $\mathbf{0.04}$ $\mathbf{0.03}$ $0.05$ $0.07$ $\mathbf{0.05}$ $\mathbf{0.01}$ $\mathbf{0.04}$ $0.04$ $\mathbf{0.25}$ $\mathbf{0.51}$ $\mathbf{0.008}$ $\mathbf{0.02}$ $\mathbf{0.2}$ $0.006$
DynoSAM Motion $\mathbf{0.04}$ $\mathbf{0.04}$ $\mathbf{0.03}$ $0.05$ $\mathbf{0.06}$ $\mathbf{0.05}$ $\mathbf{0.01}$ $\mathbf{0.04}$ $\mathbf{0.02}$ $0.64$ $\mathbf{0.51}$ $\mathbf{0.008}$ $\mathbf{0.02}$ $1.2$ $0.006$
This section explains the metrics used in our DynoSAM framework to evaluate the accuracy of pose and motion estimations for camera and objects. For camera localization accuracy, we employ conventional evaluation methods that assess the camera trajectory and motions. However, the evaluation of object pose and motion remains challenging as each object body-fixed reference frame is defined arbitrarily by each and any dataset. As explained in Sec. IV-A, different local references on the same rigid-body object will experience different local motions and have different trajectories. In addition, the information on where the object local reference frame is on each object in the ground truth is often missing from most datasets, as is the ground truth object motion. Therefore, in this section we discuss in detail the characteristics of each metric DynoSAM employs, especially those agnostic to reference frame definitions.
We evaluate the accuracy of camera pose and motion estimation using a combination of standard metrics, e.g. Absolute Pose Error (APE), Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) as defined by Sturm et al[65]. For every error metric $\mathbf{E}$ for a $\mathrm{SE}(3)$, we compute and report the root-mean-squared error (RMSE) of the translation and rotation component separately:
$\text{RMSE}(e)=\sqrt{\frac{1}{n(\mathcal{K})}\sum_{k\in\mathcal{K}}\lVert e_{k}\rVert^{2}}\text{,}$ (31)
where $e_{k}$ is a scalar error at each time-step $k$ throughout a sequence. The scalar translational error is the $L_{2}$ norm of the translation component $trans(\mathbf{E}_{k})$, and the rotational error is the angle of the rotation component $rot(\mathbf{E}_{k})$.
Given a ground truth transformation $\mathbf{M}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}, k}}\in\mathrm{SE}(3)$ and a corresponding reference estimate $\mathbf{M}_{\text{k}}\in\mathrm{SE}(3)$ at time $k\in\mathcal{K}$, the APE at this time-step is defined as:
$\mathbf{APE}_{k}=\mathbf{M}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}^{-1}\>\mathbf{M}_{k}\in\mathrm{SE}(3)\text{.}$ (32)
ATE is calculated as the root-mean-squared error (RMSE) of the translation component of the APE:
$\mathbf{ATE}=\text{RMSE}\big{(}trans\left(\mathbf{APE}_{k}\right)\big{)}\text{,}$ (33)
and is commonly used to evaluate the absolute camera trajectory error [3, 21, 35].
We use the relative pose error (RPE) metric for further pose evaluation. RPE is the RMSE of the difference in relative poses between consecutive frames, computed for translation and rotation separately:
$\mathbf{RPE}_{t}$ $=\text{RMSE}\big{(}trans\left(\mathbf{RPE}_{k}\right)\big{)}$ (34)
$\mathbf{RPE}_{r}$ $=\text{RMSE}\big{(}rot\left(\mathbf{RPE}_{k}\right)\big{)}\text{,}$
where at each time-step $k$:
@@
For evaluation of camera pose we set $\mathbf{M}=\prescript{W}{}{\mathbf{X}}_{k}$ to compute ATE and RPE. We assess object RPE by setting $\mathbf{M}=\prescript{W}{}{\mathbf{L}}_{k}$. This metric allows us to easily compare with other state-of-the-art Dynamic SLAM systems [24, 21, 35, 17]. Note that the RPE definition commonly employed here is different from the relative translation and rotation errors used by some classic visual SLAM systems to evaluate the accumulated camera pose errors over distance traveled [66, 3]. Because the poses and trajectory of an object change based on where its local reference frame is defined, which differs for each system, we do not evaluate APE or ATE for objects.
It is also vital to understand the accuracy of our estimated object motion. We evaluate the local motion of each object using the Motion Error $\mathbf{ME}$. For each estimated observed motion $\prescript{W}{k-1}{\mathbf{H}}_{k}$ we calculate the local motion using the object pose ground truth:
$\prescript{\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}}{k-1}{\mathbf{H}}_{k}=\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}^{-1}\>\prescript{W}{k-1}{\mathbf{H}}_{k}\>\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}$ (36)
and define the motion error for each time-step $k$:
$\mathbf{ME}_{k}=\prescript{L_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}}{k-1}{\mathbf{H}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}^{-1}\>\prescript{L_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}}{k-1}{\mathbf{H}}_{k}\text{,}$ (37)
As before, we report the RMSE for the translation and rotation error throughout the sequence for all objects:
$\mathbf{ME}_{t}$ $=\text{RMSE}\left(trans\left(\mathbf{ME}_{k}\right)\right)$ (38)
$\mathbf{ME}_{r}$ $=\text{RMSE}\left(rot\left(\mathbf{ME}_{k}\right)\right)\text{.}$
We include only objects that have been observed for at least $3$ consecutive frames, i.e. having at least $2$ motions in our reported results.
From Eq. (7) and our earlier definitions of motion, we highlight that RPE and ME are similar metrics, but with nuanced differences:
$\mathbf{RPE}_{k}$ $=\big{(}\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}^{-1}\>\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}\big{)}^{-1}\>\big{(}\prescript{W}{}{\mathbf{L}}_{k-1}^{-1}\>\prescript{W}{}{\mathbf{L}}_{k}\big{)}$ (39a)
$=\prescript{L_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}}{k-1}{\mathbf{H}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}^{-1}\>\prescript{L_{k-1}}{k-1}{\mathbf{H}}_{k}\ncong\mathbf{ME}_{k}\text{.}$ (39b)
The difference between Eq. (37) and the RPE component of Eq. (39a) lies in the reference frame of the estimated motions: ME uses the ground truth object frame $L_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}$ specified by the pose $\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}$ while RPE uses the estimated object frame.
Our framework includes the ME and RPE metrics as both are relevant to understanding the quality of the estimated motion and facilitate comparison with a wider range of systems. While RPE is a common criteria employed by most state-of-the-art Dynamic SLAM systems [24, 21, 17], this metric is not invariant to where the estimated object reference frame is defined. By comparison, ME is made agnostic to the estimated frame by using the ground truth object frame as its reference, which ensures that any resulting error is purely a result of error in the motion and not also in the pose. However, RPE remains a useful metric to evaluate the resulting trajectory. It should be noted that, when the Dynamic SLAM system only estimates object motions, e.g. DynoSAM Motion Estimator, the object pose $\prescript{W}{}{\mathbf{L}}_{k}$ must be computed by propagating the initial object pose $\prescript{W}{}{\mathbf{L}}_{s}$ using the estimated motions:
$\prescript{W}{}{\mathbf{L}}_{k}=\prescript{W}{k-1}{\mathbf{H}}_{k}\dots\prescript{W}{s}{\mathbf{H}}_{s+1}\>\prescript{W}{}{\mathbf{L}}_{s}\text{.}$ (40)
In this case, RPE not only accumulates the drift in motion estimation, but also sums them up throughout the object observation in the RMSE calculation.
@@
To ensure our framework performs fair and accurate evaluation, it is vital that we correctly align the camera and object frames with their respective ground truth frames. For camera pose evaluation, the estimated and ground truth states share a common and well defined reference frame, making alignment and comparison simple. During experiments we align the ground truth origin to the estimated trajectory using Umeyama’s method [67]. To facilitate a valid evaluation of RPE, we define the start of each object trajectory using the first ground truth object pose, i.e. Eq. (40) to ensure a common object frame definition. To guarantee that the estimated objects also share the same world frame, we calculate $\prescript{W}{k-1}{\mathbf{H}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}$ and $\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}$ directly from the dataset using the newly aligned ground truth odometry. This ensures that both the world and object frames are defined correctly for each sequence, facilitating a valid comparison.
VII-B Datasets
We evaluate our system on the real-world KITTI tracking dataset [66], OMD [22] and the outdoor datasets provided by ClusterSLAM [24] [^2] The datasets are provided in addition to the paper and can be found at https://huangjh-pub.github.io/page/clusterslam-dataset/ . Each dataset was selected specifically as they contain ground truth object pose information in addition to the camera pose, allowing ground truth object motion to be extracted for each object.
ClusterSLAM provides $10$ simulated dynamic sequences, $4$ of which present driving sequences similar to KITTI and include numerous instances of dynamic vehicles. These sequences are denoted as the ‘outdoor’ sequences by ClusterSLAM [24], hence we refer to them as the Outdoor Cluster dataset in this paper. These datasets provide diverse and complex scenes, including multiple objects, partial and full occlusions, and semi-rigid objects like cyclists. These challenges test DynoSAM’s ability to handle real-world scenarios. However, in most of these scenarios the objects all experience relatively constant motion.
To fully assess DynoSAM’s ability to estimate complex dynamics, we also evaluate the motion estimation results of our system using the swinging_4_unconstrained (S4U) sequence of the OMD. This sequence features four swinging cubes with unconstrained and unpredictable motion in an enclosed environment. The camera motion in this sequence also differs from KITTI and Outdoor Cluster in that it mimics hand-held camera motion with complex rotations. While the OMD contains other sequences, the S4U sequence exhibits the most challenging motion of the dataset and facilitates comparison to MVO. While MVO also utilizes the occlusion_2_unconstrained sequence to assess their occlusion handling capabilities, this specific focus is outside the scope of our work. Consequently, we do not evaluate our framework on this sequence.
We use a processed version of the KITTI tracking dataset which contains ground truth camera and object pose, as well as object motions, per frame. The Outdoor Cluster sequences use CARLA [68] as the simulation engine where ground truth object states and camera poses can be obtained directly. OMD uses a Vicon motion capture system to obtain ground truth trajectories of the camera and objects; full details of the experimental setup can be found in [22].

### v3

@@ VI-B1 Pose Evaluation
$00$ $01$ $02$ $03$ $04$ $05$ $06$ $18$ 2020 L1 L2 S1 S2 avg. S4U
ATE($\mathrm{m}$) DynaSLAM II $\underline{1.29}$ $2.31$ $0.91$ $\mathbf{0.69}$ $\mathbf{1.42}$ $\mathbf{1.34}$ $\mathbf{0.19}$ $\mathbf{1.09}$ $\underline{1.36}$ - - - - - $0.21$
VDO-SLAM $3.37$ $6.74$ $2.47$ $2.12$ $4.53$ $3.8$ $0.45$ $9.94$ $7.82$ - - - - - $0.19$
ClusterSLAM - - - - - - - - - - - - - $0.53$ -
MVO $1.53$ - - - - - - - - - - - - - $\mathbf{0.05}$
WCPE (ours) $\mathbf{0.82}$ $\mathbf{2.00}$ $\mathbf{0.73}$ $\underline{0.82}$ $\underline{2.01}$ $\underline{1.58}$ $\underline{0.31}$ $\underline{1.84}$ $\mathbf{1.26}$ $\mathbf{0.61}$ $\mathbf{0.52}$ $\mathbf{0.09}$ $\mathbf{0.13}$ $\mathbf{0.34}$ $\underline{0.11}$
WCME (ours) $\mathbf{0.82}$ $\mathbf{2.00}$ $\mathbf{0.73}$ $\underline{0.82}$ $\underline{2.01}$ $\underline{1.58}$ $\underline{0.31}$ $\underline{1.84}$ $\mathbf{1.26}$ $\mathbf{0.61}$ $\mathbf{0.52}$ $\mathbf{0.09}$ $\mathbf{0.13}$ $\mathbf{0.34}$ $\underline{0.11}$
$\text{RPE}_{r}$($\mathrm{\SIUnitSymbolDegree}$) DynaSLAM II $0.06$ $\underline{0.04}$ $\mathbf{0.02}$ $0.06$ $\mathbf{0.06}$ $\mathbf{0.03}$ $\mathbf{0.04}$ $\mathbf{0.02}$ $\mathbf{0.04}$ - - - - - -
VDO-SLAM $0.08$ $0.05$ $\underline{0.03}$ $\mathbf{0.03}$ $\mathbf{0.06}$ $\mathbf{0.03}$ $0.1$ $\underline{0.03}$ $\mathbf{0.04}$ - - - - - $0.77$
ClusterSLAM - - - - - - - - - - - - - $1.15$ -
MVO $0.19$ - - - - - - - - - - - - - $\underline{0.76}$
WCPE (ours) $\underline{0.05}$ $\mathbf{0.03}$ $\mathbf{0.02}$ $\underline{0.05}$ $\mathbf{0.06}$ $0.06$ $\underline{0.05}$ $0.04$ $\mathbf{0.04}$ $\mathbf{0.02}$ $\mathbf{0.02}$ $\mathbf{0.01}$ $\mathbf{0.02}$ $\mathbf{0.02}$ $\mathbf{0.69}$
WCME (ours) $\mathbf{0.04}$ $\mathbf{0.03}$ $\mathbf{0.02}$ $\underline{0.05}$ $\mathbf{0.06}$ $\underline{0.05}$ $\underline{0.05}$ $0.04$ $\mathbf{0.04}$ $\mathbf{0.02}$ $\mathbf{0.02}$ $\mathbf{0.01}$ $\mathbf{0.02}$ $\mathbf{0.02}$ $\mathbf{0.69}$
$\text{RPE}_{t}$($\mathrm{m}$) DynaSLAM II $\mathbf{0.04}$ $\underline{0.05}$ $\underline{0.04}$ $\mathbf{0.04}$ $\underline{0.07}$ $\underline{0.06}$ $\underline{0.02}$ $\underline{0.05}$ $0.07$ - - - - - -
VDO-SLAM $0.09$ $0.15$ $0.05$ $0.09$ $0.14$ $0.11$ $0.04$ $0.09$ $0.30$ - - - - - $0.12$
ClusterSLAM - - - - - - - - - - - - - $1.10$ -
MVO $\underline{0.07}$ - - - - - - - - - - - - - $\mathbf{0.004}$
WCPE (ours) $\mathbf{0.04}$ $\mathbf{0.04}$ $\mathbf{0.03}$ $\underline{0.05}$ $\underline{0.07}$ $\mathbf{0.05}$ $\mathbf{0.01}$ $\mathbf{0.04}$ $\underline{0.04}$ $\mathbf{0.04}$ $\mathbf{0.02}$ $\mathbf{0.01}$ $\mathbf{0.02}$ $\mathbf{0.02}$ $\underline{0.006}$
WCME (ours) $\mathbf{0.04}$ $\mathbf{0.04}$ $\mathbf{0.03}$ $\underline{0.05}$ $\mathbf{0.06}$ $\mathbf{0.05}$ $\mathbf{0.01}$ $\mathbf{0.04}$ $\mathbf{0.02}$ $\mathbf{0.04}$ $\mathbf{0.01}$ $\mathbf{0.01}$ $\mathbf{0.02}$ $\mathbf{0.02}$ $\underline{0.006}$
We use the RPE metric for camera and object pose evaluation. Unlike in some classical visual SLAM systems [Geiger13ijrr, mur2017orb] where RPE evaluates the accumulated camera pose errors over distance traveled, this RPE metric quantifies the difference in relative poses between consecutive frames and is employed by state-of-the-art Dynamic SLAM systems [Huang2019iccv, bescos2021ral, huang2020cvpr, judd2024ijrr_mvo]:
@@
We report translation and rotation error separately using (31). For camera pose, we set $\mathbf{M}=\prescript{W}{}{\mathbf{X}}_{k}$ to compute ATE and RPE. We assess object RPE by setting $\mathbf{M}=\prescript{W}{}{\mathbf{L}}_{k}$.
VI-B2 Motion Evaluation
The primary scope of our work focuses not only on poses but, more importantly, on object motion estimation. For proper motion evaluation, we must again consider how the motion changes when represented in different frames. As discussed in Sec. III-A, the local motion of an object depends directly on where the body frame is defined on the object. Since the object frame may be arbitrarily defined, there is no guarantee that the frame $\{L\}$ for any object $j$ will align between different systems and with the ground truth. Setting $\mathbf{M}=\prescript{W}{}{\mathbf{L}}_{k}$ in (33) and evaluating the object RPE, as many systems do, highlights the potential discrepancy:
$\mathbf{RPE}_{k}(\prescript{}{}{\mathbf{L}})=\prescript{L_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}}{k-1}{\mathbf{H}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}^{-1}\>\prescript{L_{k-1}}{k-1}{\mathbf{H}}_{k}\text{,}$
as this error compares the ground truth motion in the ground truth object frame with the estimated motion that is represented in the estimated object frame. Therefore, depending on the definition of $\{L\}$ relative to $\{L_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}}\}$ the reported error may be highly inconsistent.
An obvious solution is to instead directly evaluate the observed motion $\prescript{W}{}{\mathbf{H}}$. This representation is agnostic to the placement of the object frame on the rigid-body and is shared between all systems, including ground truth. However, expressing the motion error in the world frame has the unintended effect of ‘scaling’ the error’s magnitude with the objects position relative to $\{W\}$, as shown in Appendix -B. Consequently, even a small discrepancy in the actual motion can have a large error when the object is far from the origin.
To resolve both of these issues, we propose a new metric, Motion Error $\mathbf{ME}$, that evaluates all motions in the ground truth object frame as a local object frame that is common to all systems. This metric can be easily calculated from the observed motion representation using (35) or if the transform between the estimated and ground truth object frame is well defined. For each object $j$ at $k$, ME is defined as:
$\mathbf{ME}_{k}=\prescript{L_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}}{k-1}{\mathbf{H}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}^{-1}\>\prescript{L_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}}{k-1}{\mathbf{H}}_{k}$ (34)
where
$\prescript{\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}}{k-1}{\mathbf{H}}_{k}=\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}^{-1}\>\prescript{W}{k-1}{\mathbf{H}}_{k}\>\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k-1}\text{.}$ (35)
Conceptually, ME compares the ground truth motion in the ground truth frame with the estimated motion also expressed in the ground truth object frame. Equation (35) demonstrates how the estimated motion $\prescript{W}{k-1}{\mathbf{H}}_{k}$ can be expressed in the ground truth object frame using only ground truth object poses.
While both RPE and ME evaluate the accuracy in relative transformations along an objects trajectory, ME is agnostic to how each system defines the object frame. By accounting for any differences in definitions, this metric facilitates valid comparisons between systems. As most existing systems primarily evaluate RPE for objects, we report both RPE and ME metrics. However, we primarily focus on analysing ME as we believe this is a better indicator of accuracy.
As before, we report the RMSE for the translation and rotation error throughout the sequence for all objects, following (31). We report average motion error for each sequence. However, DynoSAM also outputs per-object errors for more detailed comparison if necessary. Only objects observed for at least $3$ consecutive frames are included.
VI-B3 Frame Alignment
@@
To ensure our framework performs fair and accurate evaluation, it is vital that we correctly align the world, the camera and object frames with their respective ground truth frames. During experiments we align the ground truth origin to the estimated trajectory using Umeyama’s method [umeyama1991least]. For camera pose evaluation, the estimated and ground truth states share a common and well defined reference frame, making alignment and comparison simple. To facilitate a valid evaluation of RPE, we define the start of each object trajectory using the first ground truth object pose, i.e. (19) to ensure a common object frame definition. To guarantee that the estimated objects also share the same world frame, we calculate $\prescript{W}{k-1}{\mathbf{H}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}$ and $\prescript{W}{}{\mathbf{L}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}gt}},k}$ directly from the dataset using the newly aligned ground truth odometry.
VI-C Camera Pose Error
Table I, Table II and Table III show comparisons of camera pose estimation using ATE and RPE. Our results demonstrate that in RPE, we predominantly achieve better results than all other systems. In cases where we perform worse, the difference is marginal with the maximum error difference being $0.02\text{\,}\mathrm{\SIUnitSymbolDegree}$ in rotation and $0.02\text{\,}\mathrm{m}$ in translation.
ATE is a measure of global consistency and accumulated drift. Across all datasets we demonstrate comparable and in some cases better results than other systems. Notably on the TartanAir (Shibuya) dataset, we perform better than VDO-SLAM in all cases and AirDOS in most sequences, as shown in Table II, demonstrating our robustness in environments with a large number of (non-rigid) dynamic objects. DynaSLAM II produces marginally better results over KITTI Sequence 03031818 in ATE. We believe this is due to our more conservative motion segmentation algorithm which results in relatively fewer static tracks, particularly in scenes with many objects. This introduces greater rotation bias into the visual odometry, notably on longer sequences, thus increasing the ATE.
The VIODE dataset is challenging due to the camera’s complex motions and scenarios of full occlusion by dynamic objects. Song et al. have demonstrated with their Visual-Inertial Odometry (VIO) system, DynaVINS [Song2022ral_dynavins], that, in these situations, tight IMU integration is vital for robust ego-motion estimation as no static features are available. We therefore tightly integrate the IMU into our backend using the preintegration method of Forster et al[Forster16tro] for the VIODE experiments. This highlights the flexibility of our framework to easily integrate inertial data. Table III shows DynoSAM’s superior RPE on the majority of sequences. Compared with ORB-SLAM3 [Campos2021tro] (running in IMU mode), we demonstrate superior robustness as DynoSAM completes all sequences while ORB-SLAM3 fails on three, as indicated by . While we outperform DynaVINS [Song2022ral_dynavins] in RPE, DynaVINS produces the best ATE, likely due to its comprehensive mechanism that rejects dynamic object features using IMU-informed pose priors. Consequently, DynaVINS is more robust during the transition phase of total occlusion. These situations occur in all VIODE sequences and biases DynoSAM’s eventual ATE, but do not affect our RPE accuracy on average. However, we believe that DynaVINS’s VIO mechanism is complimentary to our system and can be integrated in the future to improve overall accuracy and robustness. Running DynoSAM without the IMU results in the estimation diverging, as indicated by in Table III. We omit the numerical results due to space limitations and because it is clear from the DynaVINS paper that the IMU is needed in these challenging conditions [Song2022ral_dynavins].
Finally, both proposed DynoSAM formulations demonstrate almost identical performance. This is expected as they share a common formulation for visual odometry.
TABLE II: ATE evaluation of camera pose on TartanAir (Shibuya) [Qiu2022icra_airdos] compared against AirDOS [Qiu2022icra_airdos] (with mask) and VDO-SLAM [zhang2020vdoslam].
. TartanAir Shibuya Standing Human Road Crossing (Easy) Road Crossing (Hard) I II III IV V VI VII ATE ($\mathrm{m}$) AirDOS $0.06$ $\mathbf{0.02}$ $0.10$ $0.03$ $\mathbf{0.02}$ $0.22$ $0.56$ VDO-SLAM $0.10$ $0.61$ $0.38$ $0.39$ $0.22$ $0.24$ $0.66$ WCPE (ours) $\underline{0.03}$ $\underline{0.03}$ $\mathbf{0.02}$ $\underline{0.02}$ $\underline{0.03}$ $\underline{0.04}$ $\mathbf{0.18}$ WCME (ours) $\mathbf{0.02}$ $0.04$ $\mathbf{0.02}$ $\mathbf{0.02}$ $0.04$ $\mathbf{0.03}$ $\underline{0.19}$
Overall we demonstrate our camera pose estimation is accurate and robust in various highly dynamic environments that include rigid and non-rigid objects, full occlusion and complex camera motion. We outperform most systems on RPE and produce comparable if not better ATE results. The IMU integration on the challenging VIODE dataset further demonstrates the flexibility of our framework and the robustness of our algorithm under these challenging conditions.
VI-D Object Motion and Pose Errors
Table IV report object motion and pose errors across all sequences.  Table V summarizes the average percentage improvement of our approach over existing systems to highlight our framework’s consistent accuracy. As discussed previously, we have found comprehensive comparison challenging due to the limited number of open-source systems and sequences that each system reports on. We report both RPE and ME evaluations, but we focus on ME since we believe this metric is more accurate for evaluation between systems (Sec. VI-B2). However, existing systems do not report ME and many are closed-source. We use ‘-’ to indicate sequences where results could not be obtained.
ClusterSLAM [Huang2019iccv] only reports RPE as an accumulated average. Table V shows that our framework is upwards of $6$ times more accurate in both rotation and translation, demonstrating our superior performance.
TABLE III: ATE and RPE evaluation of camera pose on VIODE [minoda2021viode] compared against DynaVINS [Song2022ral_dynavins] and ORB-SLAM3 [Campos2021tro]. Both systems are run in stereo-inertial mode. DynoSAM is run with full IMU fusion in the backend and sequences marked with indicate failure cases (i.e. trajectory divergence) without the use of the IMU. indicates total failure cases. Mid/High labels indicate sequence difficulty and number of dynamic objects.
VIODE
City Day City Night Parking Lot
Mid High Mid High Mid High
ATE ($\mathrm{m}$) DynaVins $\mathbf{0.104}$ $\mathbf{0.150}$ $\mathbf{0.194}$ $\mathbf{0.147}$ $\mathbf{0.056}$ $\mathbf{0.065}$
ORB-SLAM3 $\underline{0.217}$ $\ast$ $1.693$ $3.006$ $\ast$ $\ast$
WCPE (ours) $2.515$ $\underline{2.128}$ $\underline{1.360}$ $\underline{2.560}$ $\underline{1.377}$ $\underline{0.764}$
WCME (ours) $2.515$ $\underline{2.128}$ $\underline{1.360}$ $\underline{2.560}$ $\underline{1.377}$ $\underline{0.764}$
$\text{RPE}_{t}$($\mathrm{m}$) DynaVins $0.024$ $0.027$ $0.019$ $\mathbf{0.023}$ $0.019$ $0.015$
WCPE (ours) $\mathbf{0.008}$ $\mathbf{0.014}$ $\mathbf{0.015}$ $0.020$ $\mathbf{0.006}$ $\mathbf{0.005}$
WCME (ours) $\mathbf{0.008}$ $\mathbf{0.014}$ $\mathbf{0.015}$ $0.020$ $\mathbf{0.006}$ $\mathbf{0.005}$
$\text{RPE}_{r}$($\mathrm{\SIUnitSymbolDegree}$) DynaVins $0.087$ $\mathbf{0.09}$ $0.102$ $\mathbf{0.096}$ $0.126$ $0.111$
WCPE (ours) $\mathbf{0.049}$ $0.105$ $\mathbf{0.070}$ $0.190$ $\mathbf{0.036}$ $\mathbf{0.040}$
WCME (ours) $\mathbf{0.049}$ $0.105$ $\mathbf{0.070}$ $0.190$ $\mathbf{0.036}$ $\mathbf{0.040}$